OpenStudy (anonymous):

The surface area of a sphere of 57pi meters. Find the radius of the sphere. Show work please.

OpenStudy (anonymous):

surface area=4 pi r^2 57=4 r^2 r^2=57/4 r=\[\sqrt{57}/2\]

OpenStudy (anonymous):

so R =3.77

OpenStudy (anonymous):

yep